OpenStudy (anonymous):

NEED HELP QUICKLY f h(x)=(fog)(x) and h(x)=5(x+1)^3, find f(x) and g(x). A. f(x)=5x^3, g(x)=(x+1)^3 B. f(x)=5x^3, g(x)=x+1 C. f(x)=x+1, g(x)=5x^3 D. f(x)=(x+1)^3, g(x)=5x^3

geerky42 (geerky42):

(fog)(x) is same as f( g(x) )

OpenStudy (anonymous):

but I dont know how to do that either because there's no g(x) and/or f(x) given

geerky42 (geerky42):

so which of the following will have \(f( g(x) ) = 5(x+1)^3\)?

geerky42 (geerky42):

well, you are given h(x) = f(g(x)), so you need to find f(x) and g(x)

geerky42 (geerky42):

hmm, you know, g(x) could be x+1 and f(x) = 5x^3 so f(g(x)) could be 5(x+1)^3

OpenStudy (anonymous):

in which case it'd be B correct?

geerky42 (geerky42):

yeah

OpenStudy (anonymous):

but it could be any combination of the equation couldn't it?

OpenStudy (anonymous):

just kidding, you got it right! Thanks!